OpenStudy (anonymous):

\[25x ^{2}-{36y^2}+900=0\] ----- I know how to work the equation, but what about the negative polynomial x^2 it leaves, how does that work?

OpenStudy (anonymous):

\[25x ^{2}-{36y^2}+900=0\]

myininaya (myininaya):

what are you trying to do @Jeremy.Kane

myininaya (myininaya):

Trying to put this in standard hyperbola form?

OpenStudy (anonymous):

convert it to a standard ellipse or parabola form, but it leaves me with the first polynomial being a negative and I can't figure out how that changes anything

myininaya (myininaya):

\[\frac{(x-h)^2}{a^2}-\frac{(y-k)^2}{b^2}=1\] This is our goal! It has to be a hyperbola if any of the conics because the x^2 and y^2 have different signs in front of it

OpenStudy (anonymous):

right, the minus sign shows that, so the negative x component doesn't change anything? I have -36^2 at the bottom, then I switch the sign to the top .. I'd get -(x+h)^2 then? that negative doesn't make a difference because it's a square?

OpenStudy (anonymous):

sorry, I should use the equation editor

myininaya (myininaya):

Or goal is really either this \[\frac{(x-h)^2}{a^2}-\frac{(y-k)^2}{b^2}=1\] or this \[\frac{-(x-h)^2}{a^2}+\frac{(y-k)^2}{b^2}=1\] This is our equation \[25x^2-36y^2=-900 \] We can tell we want one side to be 1 so that is what I try to do first the way we get one is by dividing -900/-900=1 If you divide one side, you must divide the other side by the same thing.

myininaya (myininaya):

Both of those forms I just mentioned or hyperbola forms

OpenStudy (anonymous):

your a mod, I didn't mean to give someone an answer a few minutes ago- I wasn't aware of that rule. And that is what I was thinking. the \[-(x-0)^{2} \] won't change anything because it's a square

myininaya (myininaya):

It's fine. If you are asking me if (-1)^2 is 1 or 1^2, then yes But -1^2 is not (-1)^2 which means -1^2 is not 1^2

myininaya (myininaya):

your equation is definitely not an ellipse or a parabola

myininaya (myininaya):

This is why I was saying if it is any conic it is a hyperbola

myininaya (myininaya):

http://tutorial.math.lamar.edu/Classes/Alg/Hyperbolas.aspx Here are some examples of hyperbolas. Let me know if you need more help if you return.

OpenStudy (anonymous):

im sorry, had to check with the kiddo -- it's a conic section for sure, i have to name it and find its vertices and foci - its just that negative first part that is confusing the heck out of me. I come up with

OpenStudy (anonymous):

shoot, one sec

OpenStudy (anonymous):

\[\frac{ (x-0)^{2} }{ -6^2 }\] which is the same as \[-(x-0)^{2}/{6^2}\] .. its the negative sign that is throwing me off -- really do appreciate this help

OpenStudy (anonymous):

I hope that makes sense